please answer all 3!
Find the inverse algebraically and then graph it. y-x-3
Find the inverse below. f(x) = x-4 +9 Submit
f(x) = Find the inverse below. x + 6 Submit
Answer #1

fl2)= 316-4 to Let y=f(e) = 3/574 +9. y 33-4 +9 switch te intory and solve for yo t = 3/4-4 +9 317.4 = -9 cube both sides. 4-det y = f(x) = 310th y 31 446 switch de into y and solve for y 314 = -6 cube both sides ya (-638 S- (de) = (2-6) 3. 2 4 /6 8